A bag contains 5 black balls, 4 white balls and 3 red balls. If a ball is selected randomwise, the probability that it is a black or red ball is ____ .

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Explaination: Number of black or red balls = 5 + 3 = 8
Debarati
2 years ago
No. of favourable outcomes= 5 + 3= 8
Total no. of outcomes= 5 + 4 + 3= 12
Probability of getting a black or red ball= 8/12= 2/3